identify the limiting agent if 50.0 g Ag reacts with 10.0 g S
1
Expert's answer
2020-12-10T07:50:43-0500

2Ag + S "\rightarrow" Ag2S


For one mole of sulphur there should be two moles of silver for complete reaction.

Moles of silver = 50/108 = 0.46 moles

Moles of elemental sulphur = 10/32 = 0.3125 moles

For 0.3125 moles of sulphur there must be 0.6250mole of silver.

So silver is limiting Reagent.